# Unmasking Illusions: Understanding Human Perception of Audiovisual Deepfakes

### v1

@@ Abstract
The emergence of contemporary deepfakes has attracted significant attention in machine learning research, as artificial intelligence (AI) generated synthetic media increases the incidence of misinterpretation and is difficult to distinguish from genuine content. Currently, machine learning techniques have been extensively studied for automatically detecting deepfakes. However, human perception has been less explored. Malicious deepfakes could ultimately cause public and social problems. Can we humans correctly perceive the authenticity of the content of the videos we watch? The answer is obviously uncertain; therefore, this paper aims to evaluate the human ability to discern deepfake videos through a subjective study. We present our findings by comparing human observers to five state-of-the-art audiovisual deepfake detection models. To this end, we used gamification concepts to provide 110 participants (55 native English speakers and 55 non-native English speakers) with a web-based platform where they could access a series of 40 videos (20 real and 20 fake) to determine their authenticity. Each participant performed the experiment twice with the same 40 videos in different random orders. The videos are manually selected from the FakeAVCeleb dataset. We found that all AI models performed better than humans when evaluated on the same 40 videos. The study also reveals that while deception is not impossible, humans tend to overestimate their detection capabilities. Our experimental results may help benchmark human versus machine performance, advance forensics analysis, and enable adaptive countermeasures.
@@
Deepfake detection, Video forgery detection, Audiovisual deepfake detection, Human perception, Subjective evaluation, gamification, crowdsourcing
@@ I Introduction
Deepfakes are computer-generated or manipulated multimedia content that convincingly simulates or conceals reality. The emergence of Artificial Intelligence (AI) and Deep Learning (DL) technologies has advanced manipulation tools that are capable of altering a person’s identity by swapping voice or facial features. Such fabricated multimedia content, generated or manipulated through cutting-edge AI techniques, can depict a person in situations they’ve never been in. Deepfake generation technology is a gold mine with wide applications; however, its dark side poses a significant threat to society and can have shocking consequences for unsuspecting citizens. For instance, it can cause chaos when highly realistic videos of public figures or politicians are generated to spread misinformation. Fake videos can damage someone’s reputation and emotional well-being by featuring them in explicit content without their consent. The widespread malicious and unethical use of synthetic fake content on social media platforms has caused public unrest and raised serious trust issues in the media.
@@
Exponential advances in hyper-realistic visual effects and the manipulation across various modalities, including audio, image, and video, have led to the creation of increasingly plausible content, thereby undermining the trust in the proverb “Seeing is believing” in digital media. This raises concerns that audiovisually manipulated videos become visually and aurally indistinguishable. Audiovisual deepfakes are a form of synthetic media generated via sophisticated deep learning techniques that manipulate the acoustic and visual tracks of a video, often created with the intention of deceiving people. Fig. 1 provides an illustrative example of audiovisual deepfake creation. In addition, the development of deepfake generation techniques, such as Generative Adversarial Network (GAN) [1], Autoencoder (AE), and Variational Autoencoder (VAE) [2], and their growing public accessibility have had a strong impact on visual and auditory content. This highlights the importance of caution within the research community as it can be difficult to distinguish between truth and fabrication, creating new social challenges. In this paper, we study human perception of audiovisual deepfakes. Our contribution provides a deeper understanding of human performance in detecting audiovisual deepfakes, and our analysis may be valuable for enhancing cybersecurity measures, advancing forensic analysis, and developing countermeasures.
@@
The generation [1, 3, 4, 5] and detection [5, 6, 7, 8, 9] of deepfakes have been extensively explored by the research community. While human perception of deepfake speech, images, or videos respectively has been a major focus of researchers, to date, there has been less research in the field of human perception of audiovisual deepfakes. This study is an empirical investigation designed to evaluate the ability of human participants to detect audiovisual deepfake videos. The objective of the experiment is to analyze the perceptual boundaries of audiovisual deepfakes and explore how the average human observers perceive and interpret fabricated content. Therefore, we conducted a comprehensive crowdsourced subjective test via a web-based platform, allowing human participants to watch a series of videos and identify authenticity. Evaluating human’s ability to detect deepfakes is crucial, but it is also important to understand how detection algorithms identify them. Therefore, we also tested five state-of-the-art (SOTA) AI deepfake detectors on the same set of videos. Finally, we compared the subjective test results with the test results of the AI detection models. In particular, we explored the following questions:
@@ II Related Work
For images, research on deepfake technology has shown that subtle changes in faces can generate compelling images that are difficult to detect by human observers. Therefore, numerous studies have attempted to examine human perception capabilities to identify fabricated images, and have shown that human performance is not significantly better than the random chance level. For example, the subjective study in [15] analyzed the human ability to detect forgery in digital images, and found that humans can be easily fooled and that participants not only performed poorly at identifying image forgeries, but also often doubted the authenticity of real images. [16] evaluated the human ability to identify image deepfakes of human faces, and found that people are not naturally good at detecting deepfakes and that the simple interventions tested do not help. [17] focused on a crowdsourcing-based approach to evaluate synthetic images specifically generated with GANs. The study in [18] shows that synthetically generated faces are not just highly photorealistic, they are nearly indistinguishable from real faces and are judged more trustworthy. The study in [19] explored the impact of synthetic image deepfakes on college students’ perception and understanding of the technology.
@@
There are also subjective studies analyzing human ability to distinguish between authentic and synthetic videos, and comparing human performance with AI models. The study in [20] shows that both human and AI models exhibit comparable performance, but with different errors, and that subjects with access to model predictions perform better than subjects without access to model predictions. The study in [21] shows that while human perception is very different from machine perception, both are successfully fooled by deepfakes but in different ways. Specifically, algorithms have difficulty detecting deepfake videos that are easily spotted by humans. According to [22], people cannot reliably detect deepfakes. Raising awareness and financial incentives do not improve people’s detection accuracy. People tend to mistake deepfakes as authentic videos (rather than vice versa) and overestimate their own deepfake detection abilities. Therefore, several studies have attempted to analyze the factors that influence viewers’ ability to perceive deepfake videos. The authors of [23] conducted an investigative user study and analyzed existing AI detection algorithms to uncover the unknown factors behind the detection of deepfakes. A similar study in [24] shows that two-thirds of participants were unable to accurately detect a sequence of four videos as either genuine or deepfake, and that familiarity with the subjects in the videos had a statistically significant impact on the individuals’ perception ability. Another study in [25] delved into how humans and machines perceive deepfake videos, suggesting that incorporating emotional factors and leveraging specialized visual processing may be promising strategies to enhance deepfake detection. The study in [26] explored the difference in detection performance between human observers and automated systems by comparing their abilities to identify deepfake videos in the presence of noisy channels. The interesting work in [27] examined tourists’ visit intention by watching deepfake destination videos, and found that the factors that affect the tourists’ visit intention after watching deepfake videos include information manipulation tactics, trust and media richness, while perceived deception does not influence tourists’ visit intention. The study in [28] shows that people are more likely to feel uncertain than to be misled by political deepfakes, but this resulting uncertainty, in turn, reduces trust in news on social media. The authors of [29] explored the impact of deepfake videos’ informative cues on individuals’ perceived accuracy of claims and their intentions to share non-political deepfakes, and found that high cognitive individuals were less likely to trust news and share content, and that people were more likely to believe that deepfake claims were true when informative cues were absent. The results in [30] indicate various challenges in deepfake user perception that technology developers need to address before the potential of deepfake applications can be realized for human-computer interaction.
@@
In addition to the above-mentioned human perception and cognition of single-modal synthetic (or manipulated) speech, images, and videos, there have also been a few studies exploring cross-modal human perception. For example, the authors of [31] conducted a comprehensive survey of 3,002 participants across three countries to understand humans’ ability to detect synthetic media across audio, image, and text. The results show that SOTA forgeries are nearly indistinguishable from “authentic” media, highlighting the difficulty of identifying deepfakes. The study in [32] shows that individuals are more vulnerable to, and more likely to share, video deepfakes than other forms of deepfakes (e.g., audio deepfakes and cheapfakes), and that individuals with high cognitive ability are less likely to perceive deepfakes as accurate or share them. Another experimental study in [33] compared the performance of human perception in detecting political deepfake speeches across text, audio, and video, finding that audio and visual information enables more accurate discernment than text alone.
@@ III-A Experiment Description
Figure 4: Age distribution of participants in the experiment.
@@
To assess their ability to identify the source of manipulation in situations where participants thought the video was fake, they were prompted to carefully specify which modality of the video was altered – video, audio, or both. In the case of video manipulation, participants were further asked to indicate which aspects of the visual manipulation were noticed, such as whether the lips, face, or eyes were manipulated or the audio and video were out of sync. Likewise, in the case of audio manipulation, participants were asked to choose between the audio being manipulated or being out of sync with the video. Once participants classified a video, they were provided with the correct label; in this way, participants could track their performance. It is worth noting that the correct label is either real or fake, excluding specific manipulation types in each modality. Each participant had to complete 80 rounds, i.e., judging 80 videos. Our experiment consists of two phases. Participants judged 40 videos presented in a random order in the first phase, and then again judged the same 40 videos presented in another random order in the second phase. This was to see if individuals performed consistently the second time they judged the videos. Upon completion of the 80-rounds experiment, participants were asked to indicate their level of confidence. Each participant was asked to indicate how confident they were in the accuracy of their answers. Finally, participants received feedback on the number of correctly classified samples out of the total number.
@@ III-B Participants
We excluded participants who participated in less than 80 rounds. A total of 110 participants completed the experiment, evenly distributed by gender, with half being native English speakers and the other half being non-native English speakers. Half of the participants received a forewarning note about the threat of deepfakes, while the other half did not. This allowed us to explore whether prior awareness and understanding of the risks associated with deepfakes impacted participants’ performance. The overall distribution of participants is shown in Fig. 3. The age distribution of participants is shown in Fig. 4. Additionally, we asked participants about their information technology (IT) skill level by providing four options: beginner, intermediate, good, or expert. This distribution is shown in Table I.
@@ III-C Dataset
We conducted the study using the FakeAVCeleb dataset [36], a dataset specifically designed for the task of audiovisual forgery detection. Its fake videos cover audio manipulation, video manipulation, or both, using advanced tools including including SV2TTS [37] (a transfer learning-based real-time voice cloning tool) for synthetic clone voice generation and Wav2Lip [38], Faceswap [39], and Fsgan [40] for visual manipulation. The videos in the dataset are divided into 4 categories, namely Fake-Video-Fake-Audio (FVFA), Fake-Video-Real-Audio (FVRA), Real-Video-Fake-Audio (RVFA), and Real-Video-Real-Audio (RVRA). The dataset contains videos of diversity, including four ethnic backgrounds, various geographic locations, and both genders. Fig. 5 shows some video examples.
By leveraging this dataset, we aim to delve into human behavioral patterns in identifying audiovisual deepfake videos. Incorporating audio and visual manipulations into the dataset allows us to gain valuable insights into how individuals interpret and identify fakes across modalities.
@@ III-C1 Data Selection
In our experiment, we manually selected 40 videos from the test set of the FakeAVCeleb dataset, 20 videos from the real class, and 20 videos from the fake class. In each class, 10 videos feature males, and 10 videos feature females.

### v2

@@ Abstract
The emergence of contemporary deepfakes has attracted significant attention in machine learning research, as artificial intelligence (AI) generated synthetic media increases the incidence of misinterpretation and is difficult to distinguish from genuine content. Currently, machine learning techniques have been extensively studied for automatically detecting deepfakes. However, human perception has been less explored. Malicious deepfakes could ultimately cause public and social problems. Can we humans correctly perceive the authenticity of the content of the videos we watch? The answer is obviously uncertain; therefore, this paper aims to evaluate the human ability to discern deepfake videos through a subjective study. We present our findings by comparing human observers to five state-of-the-art deepfake detection models. To this end, we used gamification concepts to provide 110 participants (55 native English speakers and 55 non-native English speakers) with a web-based platform where they could access a series of 40 videos (20 real and 20 fake) to determine their authenticity. Each participant performed the experiment twice with the same 40 videos in different random orders. The videos are manually selected from the FakeAVCeleb dataset. We found that all AI models performed significantly better than humans when evaluated on the same 40 videos. The study also reveals that while deception is not impossible, humans tend to overestimate their detection capabilities. Our experimental results may help benchmark human versus machine performance, advance forensics analysis, and enable adaptive countermeasures.
@@
Deepfake detection, Video forgery detection, Audiovisual deepfake detection, Human perception, Human behavior, Subjective evaluation, Gamification, Crowdsourcing
@@ I Introduction
Deepfakes are computer-generated or manipulated multimedia content that convincingly simulates or conceals reality. The emergence of Artificial Intelligence (AI) and Deep Learning (DL) technologies has advanced various fields [1, 2, 3, 4, 5] also manipulation tools that are capable of altering a person’s identity by swapping voice or facial features. Such fabricated multimedia content, generated or manipulated through cutting-edge AI techniques, can depict a person in situations they’ve never been in. Deepfake generation technology is a gold mine with wide applications [6]; however, its dark side poses a significant threat to society and can have shocking consequences for unsuspecting citizens. For instance, it can cause chaos when highly realistic videos of public figures or politicians are generated to spread misinformation [7, 8, 9]. Fake videos can damage someone’s reputation and emotional well-being by featuring them in explicit content without their consent [10]. The widespread malicious and unethical use of synthetic fake content on social media platforms has caused public unrest and raised serious trust issues in the media [7][9].
@@
Exponential advances in hyper-realistic visual effects and the manipulation across various modalities, including audio, image, and video, have led to the creation of increasingly plausible content, thereby undermining the trust in the proverb “Seeing is believing” in digital media. This raises concerns that audiovisually manipulated videos become visually and aurally indistinguishable. Audiovisual deepfakes are a form of synthetic media generated via sophisticated deep learning techniques that manipulate the acoustic and visual tracks of a video, often created with the intention of deceiving people. Fig. 1 provides an illustrative example of audiovisual deepfake creation. The development of neural generative models, such as Generative Adversarial Network (GAN) [11], Autoencoder (AE), Variational Autoencoder (VAE) [12], and diffusion model [13, 14, 15], as well as multimedia content generation techniques based on them [15, 16, 17, 18, 19, 20], and the growing public accessibility of these models and techniques have had a strong impact on visual and auditory content. This highlights the importance of caution within the research community as it can be difficult to distinguish between truth and fabrication, creating new social challenges. In this paper, we study human perception of audiovisual deepfakes. Our contribution provides a deeper understanding of human performance in detecting audiovisual deepfakes, and our analysis may be valuable for enhancing cybersecurity measures, advancing forensic analysis, and developing countermeasures.
Existing deepfake detection methods fall into two primary categories, unimodal [21, 22, 23] and multimodal [24, 25, 26, 27] deepfake detectors, each with specific strengths and weaknesses. Unimodal detectors, which exclusively focus on analyzing a single modality such as visual or audio, have made a substantial impact in establishing the basics of deepfake detection. Their strengths include simplicity and efficiency, focused analysis and low data requirements compared to multimodal approaches. Despite their better generalization ability, these methods are ineffective in the detection of advanced multimodal forgeries, audio deepfake detection models often miss visual manipulation, and visual deepfake detection models fail to detect audio manipulation. Multimodal detectors analyze multiple modalities such as audio and visual data streams to enhance the detection accuracy and robustness compared to unimodal approaches. However, multimodal deepfake detectors are resource-intensive or computationally expensive as they process and analyze both visual and audio tracks concurrently. Multimodal detectors can effectively detect known types of deepfakes, yet they suffer from limited generalization and may face challenges with novel or previously unseen manipulation techniques.
@@
The generation [11, 28, 29, 21] and detection [21, 24, 25, 27, 26, 30, 31] of deepfakes have been extensively explored by the research community. While human perception of deepfake speech, images, or videos respectively has been a major focus of researchers, to date, there has been less research in the field of human perception of audiovisual deepfakes. This study is an empirical investigation designed to evaluate the ability of human participants to detect audiovisual deepfake videos. The objective of the experiment is to analyze the perceptual boundaries of audiovisual deepfakes and explore how the average human observers perceive and interpret fabricated content. Therefore, we conducted a comprehensive crowdsourced subjective test via a web-based platform, allowing human participants to watch a series of videos and identify authenticity. Evaluating human’s ability to detect deepfakes is crucial, but it is also important to understand how detection algorithms identify them. Therefore, we also tested five state-of-the-art (SOTA) AI deepfake detectors on the same set of videos. Finally, we compared the subjective test results with the test results of the AI detection models. In particular, we explored the following questions:
@@ II Related Work
For images, research on deepfake technology has shown that subtle changes in faces can generate compelling images that are difficult to detect by human observers [37]. Therefore, numerous studies have attempted to examine human perception capabilities to identify fabricated images, and have shown that human performance is not significantly better than the random chance level. For example, the subjective study in [38] analyzed the human ability to detect forgery in digital images, and found that humans can be easily fooled and that participants not only performed poorly at identifying image forgeries but also often doubted the authenticity of real images. The studies in [39, 40, 41] evaluated the human ability to identify image deepfakes of human faces, and found that people are not naturally good at detecting deepfakes and that the simple interventions tested do not help. [42] focused on a crowdsourcing-based approach to evaluate synthetic images specifically generated with GANs. The study in [43] shows that synthetically generated faces are not just highly photorealistic, they are nearly indistinguishable from real faces and are judged more trustworthy. The study in [44] explored the impact of synthetic image deepfakes on college students’ perception and understanding of the technology. The study in [45] compared the ability of humans and machine learning models to detect synthetic images, with misclassification rates of 38.7% and 13%, respectively.
@@
There are also subjective studies analyzing human ability to distinguish between authentic and synthetic videos, and comparing human performance with AI models. The study in [46] shows that both human and AI models exhibit comparable performance, but with different errors, and that subjects with access to model predictions perform better than subjects without access to model predictions. The study in [47] shows that while human perception is very different from machine perception, both are successfully fooled by deepfakes but in different ways. Specifically, algorithms have difficulty detecting deepfake videos that are easily spotted by humans. According to [48], people cannot reliably detect deepfakes. Raising awareness and financial incentives do not improve people’s detection accuracy. People tend to mistake deepfakes as authentic videos (rather than vice versa) and overestimate their own deepfake detection abilities. Therefore, several studies have attempted to analyze the factors that influence viewers’ ability to perceive deepfake videos. The authors of [49] conducted an investigative user study and analyzed existing AI detection algorithms to uncover the unknown factors behind the detection of deepfakes. A similar study in [50] shows that two-thirds of participants were unable to accurately detect a sequence of four videos as either genuine or deepfake, and that familiarity with the subjects in the videos had a statistically significant impact on the individuals’ perception ability. In addition, the studies in [51, 52] investigated psychological and social factors that influence people’s ability to detect fake videos. [53] and [54] also contain interesting subjective studies, with the former investigating the role of system-generated cues, such as video quality and technical imperfections, and the latter focusing on developing and testing strategies to enhance human’s ability to identify deepfakes. Another study in [55] delved into how humans and machines perceive deepfake videos, suggesting that incorporating emotional factors and leveraging specialized visual processing may be promising strategies to enhance deepfake detection. The study in [56] explored the difference in detection performance between human observers and automated systems by comparing their abilities to identify deepfake videos in the presence of noisy channels. The interesting work in [57] examined tourists’ visit intention by watching deepfake destination videos, and found that the factors that affect the tourists’ visit intention after watching deepfake videos include information manipulation tactics, trust and media richness, while perceived deception does not influence tourists’ visit intention. The study in [7] shows that people are more likely to feel uncertain than to be misled by political deepfakes, but this resulting uncertainty, in turn, reduces trust in news on social media. The authors of [58] explored the impact of deepfake videos’ informative cues on individuals’ perceived accuracy of claims and their intentions to share non-political deepfakes, and found that high cognitive individuals were less likely to trust news and share content, and that people were more likely to believe that deepfake claims were true when informative cues were absent. The results in [59] indicate various challenges in deepfake user perception that technology developers need to address before the potential of deepfake applications can be realized for human-computer interaction.
@@
In addition to the above-mentioned human perception and cognition of single-modal synthetic (or manipulated) speech, images, and videos, there have also been a few studies exploring cross-modal human perception. For example, the authors of [60] conducted a comprehensive survey of 3,002 participants across three countries to understand humans’ ability to detect synthetic media across audio, image, and text. The results show that SOTA forgeries are nearly indistinguishable from “authentic” media, highlighting the difficulty of identifying deepfakes. Similarly, the authors of [61, 62] studied the effectiveness of human discernment in identifying AI-generated content across various media formats, including image, video, audio, and audiovisual stimuli. The authors of [62] studied the psychological and perceptual effects on people after being deceived by deepfakes. The study in [63] shows that individuals are more vulnerable to, and more likely to share, video deepfakes than other forms of deepfakes (e.g., audio deepfakes and cheapfakes [^1] Cheapfakes, also known as shallowfakes, are media content manipulated using non-AI methods, such as basic video/audio editing tools.), and that individuals with high cognitive ability are less likely to perceive deepfakes as accurate or share them. Another experimental study in [64] compared the performance of human perception in detecting political deepfake speeches across text, audio, and video, finding that audio and visual information enables more accurate discernment than text alone.
@@ III-A Experiment Description
III-A1 Participant Selection
The participants for the study were recruited through both convenience sampling and targeted recruitment, ensuring diverse representation based on age, gender, and background. There were 110 participants in the study, and participation criteria include:
  • Aged between 20 and 50 years old.
  • No history of visual or hearing impairment that might affect the ability to evaluate audiovisual content.
  • Beginner, intermediate, good, or expert level understanding of digital media or synthetic media technology.
Before participating in the experiment, participants were informed of the purpose of the study and asked to provide informed consent.
III-A2 Study Procedure
@@
To assess their ability to identify the source of manipulation in situations where participants thought the video was fake, they were prompted to carefully specify which modality of the video was altered – video, audio, or both. In the case of video manipulation, participants were further asked to indicate which aspects of the visual manipulation were noticed, such as whether the lips, face, or eyes were manipulated or the audio and video were out of sync. Likewise, in the case of audio manipulation, participants were asked to choose between the audio being manipulated or being out of sync with the video. Once participants classified a video, they were provided with the correct label; in this way, participants could track their performance. It is worth noting that the correct label is either “real” or “fake”, excluding specific manipulation types in each modality. Each participant had to complete 80 rounds, i.e., judging 80 videos. Our experiment consists of two phases. Participants judged 40 videos presented in a random order in the first phase, and then again judged the same 40 videos presented in another random order in the second phase. This was to see if individuals performed consistently the second time they judged the videos. Upon completion of the 80-rounds experiment, participants were asked to indicate their level of confidence. Each participant was asked to indicate how confident they were in the accuracy of their answers. In the end, participants received feedback on the number of correctly classified samples out of the total number.
Figure 3: Distribution of participants in the experiment.
@@ III-B Participants
We excluded participants who participated in less than 80 rounds. A total of 110 participants completed the experiment, evenly distributed by gender, with half being native English speakers and the other half being non-native English speakers. Half of the participants received a forewarning note, as shown in Fig. 4, about the threat of deepfakes, while the other half did not. This forewarning note informed participants about audiovisual manipulation, highlighted the threats posed by deepfakes, and further emphasized the need to be aware of potential visual irregularities such as inconsistencies between the speaker’s facial movements and the audio, unnatural lip movements or glitches in videos. This setting allowed us to explore whether prior awareness and understanding of the risks associated with deepfakes impacted participants’ performance.
Figure 4: Forewarning note presented to half of the participants at the beginning of the experiment.
The overall distribution of participants is shown in Fig. 3. The age distribution of participants is shown in Fig. 5. Additionally, we asked participants about their information technology (IT) skill level by providing four options: beginner, intermediate, good, or expert. This distribution is shown in Table I.
@@ III-C Dataset
We conducted the study using the FakeAVCeleb dataset [67], a dataset specifically designed for the task of audiovisual forgery detection. The dataset contains 500 videos featuring ethnically diverse celebrities from industries including music, sports, politics, and film, prioritizing diversity in ethnicity, age, geography, race, and gender. These 500 original videos were used to generate 19,500 fake video samples. The videos of 430 celebrities were used for training, and the videos of 70 celebrities were used for testing. The fake videos cover audio manipulation, video manipulation, or both, using advanced tools including SV2TTS [68] (a transfer learning-based real-time voice cloning tool) for synthetic clone voice generation and Wav2Lip [28], Faceswap [69], and Fsgan [70] for visual manipulation. The videos in the dataset are divided into 4 categories depending on whether they were acoustically or visually manipulated, namely Fake-Video-Fake-Audio (FVFA), Fake-Video-Real-Audio (FVRA), Real-Video-Fake-Audio (RVFA), and Real-Video-Real-Audio (RVRA). These detailed labels for audio and video streams facilitate training and performance evaluation of the forgery detection system. Fig. 6 shows some video examples.
We chose to use the FakeAVCeleb dataset in our study because of the following important factors: 1) video diversity, such as four ethnic backgrounds, various geographic locations, and both genders, 2) use of the latest manipulation techniques, 3) comprehensive labeling, and 4) multimodal manipulation, involving both audio and video manipulations, in line with our research focus. By leveraging this dataset, we aim to delve into human behavioral patterns in identifying audiovisual deepfake videos. Incorporating audio and visual manipulations into the dataset allows us to gain valuable insights into how individuals interpret and identify fakes across modalities.
@@ III-C1 Data Selection
In our experiment, we manually selected 40 videos from the test set of the FakeAVCeleb dataset, 20 videos from the real class, and 20 videos from the fake class. In each class, 10 videos feature males, and 10 videos feature females. No preprocessing steps were performed on the videos before they were used in the study.
@@
III-D Web Platform
The experiment was conducted using our gamification-based web platform. We leveraged the popular open-source web server nginx to engage participants in the task of identifying the authenticity of videos they watched. For backend development, we used the high-performance framework PHP Yii2 and used PHP 7.4 as the developing language. We used the PHP $shuffle()$ function based on the Mersenne Twister algorithm to ensure repeatability and randomize the videos. MongoDB, known for its flexibility and performance, was used as a NoSQL database, with MongoDB Compass providing an intuitive graphical user interface (GUI). For efficient code editing, we used Sublime Text.